GFTODVI(1) GFTODVI(1) NAME gftodvi−makeproofsheetsfromgenericfontfiles SYNOPSIS gftodvi[−overflow-label-offset=real][−verbose][gffilename] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thegftodviprogramconvertsagenericfont(gf)fileoutputby,forexample,mf(1),toadevice independent(DVI)file(thatcanthenbetypesetusingthesamesoftwarethathasalreadybeen writtenforTEX).Thecharactersinthegffilewillappearoneperpage,withlabels,titles,and annotationsasspecifiedinAppendixH(HardcopyProofs)ofTheMETAFONTbook. gftodviusesotherfontsinadditiontothemaingffile.A‘gray’fontisusedtotypesetthepixels thatactuallymakeupthecharacter.(Wewouldn’twantallthepixelstobesimplyblack,since thenlabels,keypoints,andotherinformationwouldbelost.)A‘title’fontisusedfortheinformationatthetopofthepage.A‘label’fontisusedforthelabelsonkeypointsofthefigure.A ‘slant’fontisusedtotypesetdiagonallines,whichotherwisehavetobesimulatedusinghorizontal andverticalrules.Thedefaultgray,title,andlabelfontsaregray,cmr8,andcmtt10,respectively;thereisnodefaultslantfont. Tochangethedefaultfonts,youcangivespecialcommandsinyourMETAFONTsourcefile,or youcanchangethefontsonline.Anonlinedialogensuesifyouendthegffilenamewitha‘/’. Forexample, gftodvicmr10.300gf/ Specialfontsubstitution:grayfontblack OK;anymore?grayfontarea/home/art/don/ OK;anymore?slantfont/home/fonts/slantimagen6 OK;anymore?<RET> willuse/home/art/don/blackasthe‘gray’fontand/home/fonts/slantimagen6asthe‘slant’font (thisnameindicatesafontforlineswithslope1/6attheresolutionofanImagenprinter). Thegffilenameonthecommandlinemustbecomplete.(Theprogrampromptsyouforitifyou don’tgiveit.)Becausetheresolutionispartoftheextension,itwouldnotmakesensetoappend adefaultextensionasisdonewithTEXorDVI-readingsoftware.Theoutputfilenamedefaultsto thesamerootasthegffile,withthedviextensionadded.Forexample,theinputfile cmr10.2602gfwouldbecomecmr10.dvi. OPTIONS Theargumentto−overflow-label-offsetspecifiesthedistancefromtherightedgeofthecharac-terboundingboxatwhichtheoverflowequations(ifany)aretypeset.ThevalueisgiveninTEX points.Thedefaultisalittleovertwoinches. Withoutthe−verboseoption,gftodvioperatessilently.Withit,abannerandprogressreport areprintedonstdout. ENVIRONMENT gftodvilooksforgffilenameusingtheenvironmentvariableGFFONTS.Ifthatisnotset,it usesthevariableTEXFONTS.Ifthatisnotset,itusesthesystemdefault. Seetex(1)forthedetailsofthesearching. FILES {gray.tfm,...} Thedefaultfonts. {gray.mf,...} TheMETAFONTsources. Web2C7.32x 14December1993 1 GFTODVI(1) GFTODVI(1) SEEALSO tex(1),mf(1). DonaldE.Knuth,TheMETAFONTbook(VolumeCofComputersandTypesetting),Addison-Wes-ley,1986,ISBN0-201-13445-4. DonaldE.Knuthetal.,METAFONTware. AUTHORS DonaldE.Knuthwrotetheprogram.ItwaspublishedaspartoftheMETAFONTwaretechnical report,availablefromtheTEXUsersGroup.PaulRichardsportedittoUnix. Web2C7.32x 14December1993 2